if a bike goes around a circle that has a radius of 501 meters with an acceleration of 1.00 m/s/s what was the bike’s speed


1
Expert's answer
2021-02-03T16:12:32-0500

The acceleraion of the circular motion is given as follows:


"a = \dfrac{v^2}{R}"

where "v" is the linear speed, and "R = 501m". Is the radius of the circle. Since "a = 1m\/s^2", expressing "v" obtain:


"v = \sqrt{aR}\\
v = \sqrt{1\cdot 501} \approx 22.4m/s"

Answer. 22.4 m/s.